OpenStudy (purple_pink):

Which equation is represented by the table? x y (x, y) 0 –1 (0, –1) 1 4 (1, 4) 2 9 (2, 9) A. y = 3x – 4 B. y = 5x – 1 C. y = 3x + 3 D. y = –3x + 3

OpenStudy (purple_pink):

@Ilovecake @confluxepic @CloverRacer @Carolinapatherlover @mathstudent55 @Hotchellerae21

OpenStudy (purple_pink):

help me please

OpenStudy (purple_pink):

@confluxepic

OpenStudy (purple_pink):

help me

OpenStudy (ilovecake):

D.

OpenStudy (purple_pink):

thanks can you help me with 1 more

OpenStudy (ilovecake):

okay

OpenStudy (purple_pink):

but first could you explain or show in a way how you got that answer

OpenStudy (purple_pink):

please

OpenStudy (purple_pink):

nevermind um i will post the other question

OpenStudy (purple_pink):

http://static.k12.com/calms_media/media/713500_714000/713812/3/e7f40ad4904c683155cb05edc1e8e4059581a0a5/74440_question.jpg Which equation is shown on the graph? A. y = x + 2 B. y = 2x C. y = x – 2 D. y = 2x + 2

OpenStudy (ilovecake):

I am taking my own test and i was coming to ask my questions. So I don't really have time. But, I am 100% sure it is D.

OpenStudy (purple_pink):

help pliz

OpenStudy (ilovecake):

That one is A.

OpenStudy (purple_pink):

If x can be any number, how many solutions are there for the equation? y = –2x + 4 A. There are many solutions. B. There are two solutions. C. There is only one solution. D. There is no solution.

OpenStudy (purple_pink):

last one

OpenStudy (purple_pink):

@Ilovecake

OpenStudy (purple_pink):

@CarinaKikiDemalleo can you pliz help me

OpenStudy (anonymous):

What u need hun?

OpenStudy (anonymous):

I Think the answer is A, but Not too sure

OpenStudy (ilovecake):

Sorry I am back. I was just doing something. But Yeah @CarinaKikiDemalleo is right. It is A.

OpenStudy (anonymous):

Yay I'm right :D

OpenStudy (purple_pink):

yea it is A thanks a lot